The consecutive integers below add up to 447.
x-2
x-1
x
x+1
x+2
x+3
What is the value of x

Answer by checkley77(12844)

x-2+x-1+x+x+1+x+2+x+3=447
6X+3=447
6X=447-3
6X=444
X=444/6
X=74 ANS.